OpenStudy (anonymous):

write an equation of the line containing the given point and perpendicular to the given line (8,-8); 4x+9y=7 The equation of the line is y = ?

OpenStudy (anonymous):

9x-4y=c to determine c plug x=8,y=-8 so 72+32=c=104 so we have 9x-4y=104.. i assume u know that equation of a line perpendicular to ax+by+c=0 is bx-ay+k=0 where c and k are different constants.. so our answer is y=(1/4)(9x-104)

OpenStudy (anonymous):

jazi1 do u get tat? :)

OpenStudy (nowhereman):

jazi1: could you please try and _understand_ how it is done, and not let the others do all these problems, which are merely the same with different numbers!

OpenStudy (anonymous):

nowhereman seems like you are the only one complaining

OpenStudy (anonymous):

saubhik I got it but I don't understand it

OpenStudy (anonymous):

i think nowhereman is right:) but u shud understand what i did...or ask what u didnt get in this process instead of asking similar problems...:) happy to help... :)

OpenStudy (anonymous):

line 1: 4x + 9y = 7 It should help you to rearrange this equation with y on one side in order to resemble the standard y = mx + b format. It's critical because you need to identify the slope (m) in order to recreate a parallel or perpendicular line. In our case: 4x + 9y = 7 9y = 7 - 4x y = 7/9 - (4/9)x y = (-4/9)x + 7/9 Can you identify which is the slope (m) in this equation?

OpenStudy (anonymous):

ok...jazi1 see if we have perpendicular lines we have product of their slopes as -1. so u can tell that ax+by+c=0 and bx-ay+k=0 are 2 perpendicular lines..i used this idea..do u get this? :)

OpenStudy (anonymous):

saubhik got it

OpenStudy (anonymous):

so ur given equation is 4x+9y-7=0..so u know that the perpendicular to it has to be of the form 9x-4y+c=0 where we have to determine the constant c..do u get this? :)

OpenStudy (anonymous):

For finding c we just plug the given coordinates! thats it..nice and easy! :)

OpenStudy (anonymous):

understood jazi1?

OpenStudy (anonymous):

-4x+7 y=--------- = -4x/9 +7/9, here slope m= -4/9 9 now at point(8,-8), from y-y1=m(x-x1) y-(-8)= -4/9 (x-8) y+8 = -4x/9 +32/9 y= -4x/9 +32/9 - 8 y=-4x/9 - 40/9 EQ of the line at point(8,-8)

OpenStudy (anonymous):

Mark I do thank you but the correct ans is 9/4x - 26

OpenStudy (anonymous):

ok good try to review the process i might have missed something there..lol just in a hurry lol

OpenStudy (anonymous):

did you see it jazi? lol sorry iwas just in a hurry to give it to you lol...

OpenStudy (anonymous):

ok i saw it m=9/4 y-y1=m(x-x1) y-(-8)=9/4(x-8) y+8 =9x/4 - 72/4 y = 9x/4 -18 -8 y= 9x/4 -26 answer

OpenStudy (anonymous):

y= 9x/4 -26 this is the EQ of perpendicular line to it

OpenStudy (anonymous):

Ok thanks mark for being so helpful

OpenStudy (anonymous):

wc just understand the process of geting the Eq of the line and the EQ of the line Perpendicular to it..good luck Jazil

OpenStudy (anonymous):

saubnik did the other way of geting the line perpendicular to it which i think a kind of advance for the beginers but it is very good techniques and good process too

OpenStudy (anonymous):

if a^3 = b^3 + c^3 + d^3 , then least value of a is ?